OpenStudy (anonymous):

factor 4m^2-5n^2 i think i got the right answer for this but im not shore

jhonyy9 (jhonyy9):

us a^2 -b^2 =(a-b)(a+b) - so where a=2m and b=nsqrt5 hope this will help you